Ab Initio Calculation of U(BH_{4})_{4}

\underline{Zhiyong Zhang}, Russell M. Pitzer, Department of Chemical Physics, The Ohio State University, Columbus, OH43210.

SCF calculations with relativistic core potential and double zeta Gaussian basis set, followed by spin-orbit CI calculations, were carried out on U(BH_{4})_{4}. The ground state is found to be E state in double group notation, in agreement with experiments. The low lying f− >f transitions are calculated and results are compared with experiments and ligand field theory assignments.